A research group conducted an extensive survey of 2948 wage and salaried workers on issues ranging from relationships with their bosses to household chores. The data were gathered through hour-long telephone interviews with a nationally representative sample. In response to the question, "What does success mean to you?" 1541 responded, "Personal satisfaction from doing a good job." Let p be the population proportion of all wage and salaried workers who would respond the same way to the stated question. Find a 90% confidence interval for p. (Round your answers to three decimal places.)

lower limit    
upper limit    
Answer #1

Solution :

Given that,

n = 2948

x =1541

Point estimate = sample proportion = = x / n = 1541 /2948 =0.523

1 - = 1 -0.523=0.477

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 1.645 (((0.523 *0.477) / 2948)

= 0.015

A 90% confidence interval for population proportion p is ,

- E < p < + E

0.523 - 0.015 < p <0.523 +   0.015

0.508< p < 0.538

lower limit=0.508

upper limit=0.538